Hamming & SECDED Code

(a) For data, using 3 Hamming code parity bits determine the maximum number of data bits that can be protected.

(b) A SECDED encoded character has been retrieved, with the hexadecimal value of 409(base16). You may assume that the SECDED parity is even.

1. Was there an error in transmission? Explain your answer.

2. If there was an error, either correct it (reporting the corrected ASCII character) or explain why it could not be corrected (Show your Hamming/SECDED table).
Answer #1

(a) For data, using 3 Hamming code parity bits determine the maximum number of data bits that can be protected.

A byte of data: 10011010
Create the data word, leaving spaces for the parity bits: _ _ 1 _ 0 0 1 _ 1 0 1 0
Calculate the parity for each parity bit (a ? represents the bit position being set):

  • Position 1 checks bits 1,3,5,7,9,11:
    ? _ 1 _ 0 0 1 _ 1 0 1 0. Even parity so set position 1 to a 0: 0 _ 1 _ 0 0 1 _ 1 0 1 0
  • Position 2 checks bits 2,3,6,7,10,11:
    0 ? 1 _ 0 0 1 _ 1 0 1 0. Odd parity so set position 2 to a 1: 0 1 1 _ 0 0 1 _ 1 0 1 0
  • Position 4 checks bits 4,5,6,7,12:
    0 1 1 ? 0 0 1 _ 1 0 1 0. Odd parity so set position 4 to a 1: 0 1 1 1 0 0 1 _ 1 0 1 0
  • Position 8 checks bits 8,9,10,11,12:
    0 1 1 1 0 0 1 ? 1 0 1 0. Even parity so set position 8 to a 0: 0 1 1 1 0 0 1 0 1 0 1 0
  • Code word: 011100101010.

(b) A SECDED encoded character has been retrieved, with the hexadecimal value of 409(base16). You may assume that the SECDED parity is even.

Length of Data Sequence Number of Parity Bits Length of CodeWord
12 5 17


Parity Scheme

P0 = D0 + D1 + D3 + D4 + D6 + D8 + D10 + D11
P1 = D0 + D2 + D3 + D5 + D6 + D9 + D10
P2 = D1 + D2 + D3 + D7 + D8 + D9 + D10
P3 = D4 + D5 + D6 + D7 + D8 + D9 + D10
P4 = D11

Syndrome Matrix

D11 D10 D9 D8 D7 D6 D5 D4 D3 D2 D1 D0 P4 P3 P2 P1 P0
1 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0
0 1 1 1 1 1 1 1 0 0 0 0 0 1 0 0 0
0 1 1 1 1 0 0 0 1 1 1 0 0 0 1 0 0
0 1 1 0 0 1 1 0 1 1 0 1 0 0 0 1 0
1 1 0 1 0 1 0 1 1 0 1 1 0 0 0 0 1

1)The checksum matrix shows that the received codeword contains no errors.